# Expanding the Boundaries of Vision Prior Knowledge in Multi-modal Large Language Models

### v1

@@ 3.4 Vision Prior Remediation
Figure 7: Improvements of our training framework compared to Zero-shot. With the CLIP encoder, all three models show significant performance gains across different $Rank_{e}$ entities, particularly for those with low vision priors.
In previous sections, we established that MLLM performance correlates positively with vision prior knowledge, and that end-to-end Finetuning yields insufficient. Based on these findings, we propose VisPRE, a training framework that injects entity-related prior knowledge at vision encoder level to enhance MLLM performance. The specific process of our training framework is illustrated in Fig. 6, which comprises two key stages:
@@
DFN ViT-H-14 Vicuna-7B 59.07 58.70 63.33 64.97 62.90 65.56
Llama3.1-8B 38.70 39.84 45.08 46.99 45.69 46.99
Qwen2.5-7B 40.45 38.10 43.33 44.66 46.76 45.40
@@
  • Remedy Encoder: We first generate image captions from entity text to form (image, caption) pairs, then finetune the vision encoder alongside the text encoder using contrastive loss.
  • Instruction Tuning: We subsequently integrate the finetuned encoder into the MLLM architecture and finetune the entire model using Knowledge-tuning data.
@@
To systematically evaluate our proposed framework, we establish several baselines: Zero-shot, Perception-tuning, and Knowledge-tuning from Section 3.2. Additionally, we include Knowledge-tuning* and Mix-tuning*, where the asterisk (*) denotes unfreezing the vision encoder parameters during fine-tuning. Mix-tuning represents a combination of Knowledge-tuning and Perception-tuning data. The evaluation results are presented in Tab. 2.
@@
Finding 3: Remediating prior knowledge at vision encoder level is effective. First, Perception-tuning demonstrates only marginal improvements over Zero-shot performance, and in some configurations, performance even degrades. Knowledge-tuning improvements remain limited. This evidence suggests that fine-tuning with frozen encoder parameters fails to effectively enhance MLLM capabilities. Furthermore, Knowledge-tuning* shows modest performance gains over standard Knowledge-tuning. Finally, our VisPRE framework outperforms all baseline approaches, demonstrating substantial improvement and achieving superior results in six out of nine model combinations. These findings indicate that enhancing the encoder’s prior knowledge substantially expands the capability boundaries of MLLMs.
4 In-depth Analysis
In this section, we conduct an in-depth analysis of the reason behind our method’s effectiveness. In Section 4.1, We analyzed two possible causes: gradient diminishment and loss misalignment. In Section 4.2, we provide a case study.
4.1 Unfreezing vs. Contrastive Learning
Notably, although Knowledge-tuning* adjusted the encoder’s parameters during fine-tuning, it did not show a performance advantage over Knowledge-tuning and was outperformed by our method, which first finetune the encoder with contrastive loss before fine-tuning the MLLM. There are two possible explanations:
Vision Encoder LLM Freeze LLM All parameters
SigLIP ViT-SO-14 Vicuna-7B 29.40 52.04
Llama3.1-8B 33.47 36.58
Qwen2.5-7B 48.31 33.89
Table 3: Comparison of Freeze LLM and All parameters Strategy
Gradient Diminishment:
The gradients propagating to the encoder layer may diminish significantly, limiting their effectiveness in adjusting the encoder’s parameters. To validate the impact of gradient propagation, we conducted experiments using two distinct fine-tuning configurations. In the first configuration Freeze LLM, we kept the language model parameters fixed while finetuning only the adapter and vision encoder. In the second configuration All Parameters, we finetune all model components, including the language model, adapter, and vision encoder. Both configurations were trained on a specially constructed Perception-tuning dataset. Details of the dataset are shown in Appendix A.
The experimental results across three model combinations are presented in Tab. 3. Models fine-tuned with All Parameters consistently demonstrated superior performance compared to their Freeze LLM counterparts, showing significant improvements in four model combinations. Notably, even though Freeze LLM ensures strong gradient flow to the vision encoder due to fewer trainable layers, it still underperformed relative to All Parameters. This suggests that diminished gradients are not the primary factor limiting the effectiveness of fine-tuning with frozen parameters.
Vision Encoder LLM Genrative Contrastive
SigLIP ViT-SO-14 Vicuna-7B 52.04 50.94
Llama3.1-8B 36.58 36.99
Qwen2.5-7B 33.89 34.04
Table 4: Comparison of Generative and Contrastive Strategy
Loss Misalignment:
The language model’s predict-next-token loss may not effectively fine-tune the vision encoder, whereas contrastive loss does. To validate the Loss Misalignment hypothesis, we conducted experiments using two datasets: Generative and Contrastive. The Generative dataset is identical to Perception-tuning and serves to evaluate the effectiveness of predict-next-token loss. For the Contrastive dataset, we drew inspiration from CLIP’s contrastive learning methodology to create multiple-choice questions matching Perception-tuning’s size. In this dataset, models must select the correct entity text from several candidates that best matches the given image, allowing us to assess the effectiveness of contrastive loss. During fine-tuning on both datasets, we unfroze the encoder parameters. The results of these experiments are presented in Tab. 4. Our findings reveal that models trained with the Contrastive approach consistently outperform those trained with the Generative approach, indicating that contrastive training is more effective for knowledge acquisition in CLIP. These results provide support for our Loss Misalignment hypothesis.
4.2 Case Study
@@
Here we present an illustrative example. As shown in the upper left of Fig. 8, we input an image of the Portuguese Synagogue with the entity-related question: “Where were this synagogue’s books sent in 1979?”. For (1) LLM: The MLLM correctly answers when receiving only the textual description This is Portuguese Synagogue” instead of the actual image, indicating the LLM component possesses knowledge about this entity. For (2) MLLM (Original): With image input, the MLLM fails to answer correctly. We calculated this entity’s $Rank_{e}$ as 516, indicating low prior knowledge in the visual encoder. (3) MLLM (SFT), despite end-to-end fine-tuning, still fails since the visual encoder’s prior knowledge remains unchanged. Our training framework VisPRE first injects prior knowledge into the visual encoder, elevating the entity’s $Rank_{e}$ to 10, then conducting end-to-end fine-tuning. Consequently, (4) Ours* overcomes the visual encoder’s limitations and correctly answers the question.
@@ 5 Related Works
MLLMs incorporate visual features into language models, enabling them to perform a wide range of visual tasks. The current MLLM implementations can be primarily classified into two categories. (1) Monolithic MLLMs. Tokenizing different modal inputs uniformly and train the model from scratch[34, 5, 9, 41], which is computationally expensive. (2) Modular MLLMs. Utilizing pre-trained vision-language models (e.g., CLIP[33], SigLIP[40], DINOv2[31]) to obtain visual representations of images, and then train MLLMs through cross-modal data, aligning the visual features provided by vision encder to language model’s embedding space. This method is more data-efficient and widely used by open-source MLLMs (e.g., Flamingo[2], BLIP2[20], LLaVA[26], Qwen-VL[4]). Our work focuses on modular multimodal models. While most works treat modular MLLM as a unified system, our research focuses on the impact of vision encoder part on the language model part.
@@
With the increasing adoption of vision encoder integrated MLLMs, there are researches focusing on the relationship between vision encoder and MLLM performance. Tong et al. [37] found that CLIP[33] and corresponding MLLMs have similar performance trends across various visual modalities, indicating that CLIP features are the main cause of MLLM deficiencies in these visual modes, and addressed these deficiencies by introducing DINOv2[31] features. The recent Yang et al. [39] proposed cross-modal alignment and correspondence metrics to measure vision encoder performance, fitting a binary quadratic polynomial that accurately predicts the performance of MLLMs using that vision encoder. Different from previous works, our work provides a new perspective, showing that from a entity-related prior knowledge standpoint, CLIP’s knowledge is positive correlated with the corresponding MLLM’s performance.
@@ 6 Conclusion
In this paper, we introduce $Rank_{e}$ to quantify prior knowledge in vision encoder. We find that MLLM’s performance positively correlated with prior knowledge of vision encoder, and end-to-end finetuning MLLM yields insufficient on improving low prior entity performance. To address this issue, we propose VisPRE training framework which enhances MLLM’s performance by increasing the prior knowledge within the vision encoder. Our study demonstrates a novel pathway for enhancing MLLM performance, offering substantial value for applications involving uncommon entities.
A Datasets
Here we describe the detailed construction process of our dataset. Based on Encyclopedia-VQA[28], we constructed Knowledge-tuning and Perception-tuning datasets for each encoder-language model combination to validate Finding 2. Moreover, we developed Generative and Contrastive dataset to examine the Loss Misalignment hypothesis proposed in 4.2.
@@ A.1 Preprocess
Question Filtering. First, we focus on improving the parts where MLLM’s capabilities are limited by the vision encoder. Therefore, we only retained questions that could be answered by the corresponding LLM when prompted with ”This is {entity_name}” instead of the actual image. Next, to ensure that there were no duplicate or similar questions for the same entity across training and test sets, we deduplicated the dataset based on (entity_name, answer) pairs. Finally, we only retained entities with three or more corresponding questions to ensure sufficient questions for dividing into training and validation sets.
Prior Calculation. We calculated $Rank_{e}$ for all entities in the filtered dataset. We examined the distribution of $Rank_{e}$ calculated using different types of encoders (CLIP[33], SigLIP[40], DFN[15]) across the dataset, as shown in Fig. 10. We found significant variations in $Rank_{e}$ distributions among different encoders. CLIP’s $Rank_{e}$ values were mostly concentrated in the range of $Rank_{e}<400$, with entity counts increasing as $Rank_{e}$ decreased; In contrast, SigLIP’s $Rank_{e}$ distribution is more uniform, with at least 10 entities present across most $Rank_{e}$ intervals; DFN’s $Rank_{e}$ distribution was similar to CLIP’s, with most values concentrated in the range of $Rank_{e}<400$.
Entity Sampling. For SigLIP, we divided $Rank_{e}$ into intervals of size 1000 and sampled 10 entities from each interval. For CLIP and DFN, using the same sampling strategy as SigLIP would result in insufficient sampling of entities in dense intervals, making it difficult to distinguish different levels of prior knowledge in these regions. Therefore, we adopted a sampling method that approximates the original distributions of CLIP and DFN. We sampled 10 entities from intervals of $20<Rank_{e}<=$ 2, $42<Rank_{e}<=$ 4, $84<Rank_{e}<=$ 8, …, $1024512<Rank_{e}<=$ 1024, $Rank_{e}>1024$, ensuring that the sampled distribution approximates the original distribution while retaining all entities with low prior knowledge to reflect the relationship between entity prior knowledge and model performance. Finally, we retained the questions corresponding to the sampled entities and divided the dataset into training and test sets, with statistical information shown in Tab. 1.
@@ A.2 Construction
For Knowledge-tuning dataset, we use the original question and answer from the Encyclopedia-VQA dataset. For Perception-tuning dataset, we replace the original question in the Knowledge-tuning dataset with cognitive question like ”What is this image about?” and substitute the answers with the entity text corresponding to the image. Examples of Knowledge-tuning and Perception-tuning datasets are shown in Fig. 9.
For the Generative dataset, we use the same question and answer as in the Perception-tuning dataset but adopt a different prompt template to align with the template used in Contrasting dataset. For Contrasive dataset, instead of directly asking the model to identify the entity in the image, we provide two possible options and ask the model to choose one. We randomly adopt two types of prompts: the first is ”What is this image about?”, and the second is ”Which of the following is NOT in the image?”. For the latter, the model needs to identify the object not present in the image. The prompt templates for the Generative and Contrasive datasets are shown in Fig. 11 and Fig. 12, respectively.
B Evaluation Settings
We employ Llama-3.1-70B[14] to evaluate the accuracy of MLLM’s responses to VQA questions. Specifically, we provide Llama-3.1-70B with the question, entity name (wikipedia_title in prompt), ground truth answer, and MLLM’s response. The model outputs true to indicate a correct answer and false to indicate an incorrect answer. The prompt template is shown in Fig. 13, with the few_shot_examples shown in Fig. 14.
C More Cases
@@
Prompt for Generative dataset You are given an image. Identify what the image is about with one word or short phrase. Example: [Here is an image of Baron Hotel.] Question: What is this image about? Answer: Baron Hotel Now we begin: {entity_image} Question: What is this image about? Answer:
Figure 11: Prompt for Generative dataset. The placeholder entity_image represents the input image
Prompt for Contrasive dataset You are given an image and two options. Your task is to answer based on one of the following two types of questions: Identify what the image is about by selecting one word or short phrase from the options. Identify which of the two options is NOT present in the image. Example: [Here is an image of Baron Hotel.] Question: What is this image about? Options: (A) Markenfield Hall (B) Baron Hotel Answer: Baron Hotel [Here is an image of Baron Hotel.] Question: Which of the following is NOT in the image? Options: (A) Markenfield Hall (B) Baron Hotel Answer: Markenfield Hall Now we begin: {entity_image} Question: {question} Options: (A) {entityA} (B) {entityB} Answer:
Figure 12: Prompt for Contrasive dataset. The placeholder entity_image represents the input image, while question is selected from either ”What is this image about?” or ”Which of the following is NOT in the image?”. entityA and entityB include one correct entity corresponding to the image and one incorrect entity randomly sampled from the dataset.
Prompt for Llama-3.1 evaluation You are an expert evaluator tasked with assessing the correctness of model predictions. Your job is to determine if a given prediction is correct based on the provided information. Follow these strict guidelines: 1. You will be given four pieces of information: - Question: The original question asked - Wikipedia_title: The title of the Wikipedia article that corresponds to the knowledge base for the question - Answer: The correct answer(s) to the question, possibly including multiple candidates separated by ”|||| - Prediction: The model’s prediction to be evaluated 2. Understand that the question is specifically about the entity described in the Wikipedia_title. 3. Compare the prediction to the answer(s), taking into account the context of the question and the Wikipedia_title. 4. Apply these strict criteria: - The prediction must be accurate and specific. - If there are multiple candidate answers separated by ”||||”, the prediction must match at least one of them to be considered true. - For numerical answers, the prediction must be within 10% of at least one correct answer to be considered true. - For categorical or descriptive answers, the prediction must match the key concepts or categories in at least one of the provided answers. - Partial or vague answers that don’t fully capture the specificity of any correct answer should be considered false. - Pay close attention to units, specificity, and context provided in the question, Wikipedia_title, and answer(s). 5. Your response must be exactly one word: - Output ”true” if the prediction meets all the criteria for correctness. - Output ”false” if the prediction fails to meet any of the criteria. 6. Do not provide any explanations or additional comments. {few_shot_examples} Remember, your task is to evaluate the correctness of the prediction based on all the information provided. Be strict in your assessment, but consider all given correct answers. Respond only with ”true” or ”false”. Question: {question} Wikipedia_title: {wikipedia_title} Answer: {answer} Prediction: {prediction} Evaluation:

### v2

@@ 3.4 Vision Prior Remediation
Figure 7: VisPRE outperforms on all Ranke levels. The figure shows performance gains over Zero-shot: With the CLIP encoder, all three models demonstrate improvements across different $Rank_{e}$ entities, especially for low-prior (high $Rank_{e}$) entities.
In previous sections, we established that MLLM performance correlates positively with vision prior knowledge, and that end-to-end fine-tuning yields insufficient. Based on these findings, we propose VisPRE, a training framework that injects entity-related prior knowledge at the vision encoder level to enhance MLLM performance. The specific process of our training framework is illustrated in Fig. 6, which comprises two key stages:
@@
DFN ViT-H-14 Vicuna-7B 59.07 58.70 63.33 64.97 62.90 66.85
Llama3.1-8B 38.70 39.84 45.08 46.99 45.69 48.29
Qwen2.5-7B 40.45 38.10 43.33 44.66 46.76 43.69
@@
  • Remedy Encoder: We first reformat the Perception-tuning data into (image, entity_name) pairs, and then fine-tune the vision encoder alongside the text encoder using contrastive loss. This stage enhances the encoder’s prior knowledge of entities present in the Perception-tuning data.
  • Instruction Tuning: We incorporate the fine-tuned encoder into the MLLM architecture and perform end-to-end fine-tuning of the entire model using Knowledge-tuning data. This stage aligns the trained vision encoder with the base LLM and stimulates the model’s knowledge of entities.
@@
To systematically evaluate VisPRE, we establish several baselines: Zero-shot, Perception-tuning, and Knowledge-tuning from Section 3.2. Additionally, we include Knowledge-tuning* and Mix-tuning*, where the asterisk (*) denotes unfreezing the vision encoder parameters during fine-tuning. Mix-tuning represents a combination of Knowledge-tuning and Perception-tuning data. The evaluation results are presented in Table 2.
@@
Finding 3: Remediating prior knowledge at the vision encoder level is effective. Perception-tuning shows only marginal improvements over Zero-shot performance, occasionally even degrading results. Knowledge-tuning yields limited gains, with Knowledge-tuning* showing only modest improvement over standard Knowledge-tuning. Mix* doesn’t exceed Knowledge* performance. In contrast, our VisPRE framework outperforms all baselines, achieving superior results in six of nine model combinations. As shown in Fig. 7, VisPRE improves MLLM performance across all Ranke entities, particularly those with low vision priors, demonstrating clear advantages over alternative tuning approaches in Fig. 5. These results confirm that enhancing encoder prior knowledge substantially expands MLLM capabilities.
4 Case Study
@@
Here we present an illustrative example. As shown in the upper left of Fig. 8, we input an image of the Portuguese Synagogue with the entity-related question: “Where were this synagogue’s books sent in 1979?”. For (1) LLM: The MLLM correctly answers when receiving only the textual description “This is Portuguese Synagogue” instead of the actual image, indicating the LLM component possesses knowledge about this entity. For (2) MLLM (Original): With image input, the MLLM fails to answer correctly. We calculated this entity’s $Rank_{e}$ as 516, indicating low prior knowledge in the visual encoder. (3) MLLM (SFT), despite end-to-end fine-tuning, still fails since the visual encoder’s prior knowledge remains unchanged. Our training framework, VisPRE, first injects prior knowledge into the visual encoder, elevating the entity’s $Rank_{e}$ to 10, then conducts end-to-end fine-tuning. Consequently, (4) Ours* overcomes the visual encoder’s limitations and correctly answers the question.
@@ 5 Related Works
MLLMs incorporate visual features into language models, enabling them to perform a wide range of visual tasks. The current MLLM implementations can be classified into two categories. (1) Monolithic MLLMs. Tokenizing different modal inputs uniformly and training the model from scratch (Team, 2024a; Bavishi et al., 2023; Chen et al., 2024b; Zhan et al., 2024), which is computationally expensive. (2) Modular MLLMs. Utilizing pre-trained vision-language models (e.g., CLIP Radford et al. (2021), SigLIP Zhai et al. (2023), DINOv2 Oquab et al. (2023)) to obtain visual representations of images, and then train MLLMs through cross-modal data, aligning the visual features provided by vision encoder to language model’s embedding space. This method is more data-efficient and widely used by open-source MLLMs (e.g., Flamingo Alayrac et al. (2022), BLIP2 Li et al. (2023b), LLaVA Liu et al. (2024b), Qwen-VL Bai et al. (2023), InternVL2 Chen et al. (2024b)). Our work focuses on modular multimodal models. While most works treat modular MLLM as a unified system, our research focuses on the impact of vision encoder part on the language model part.
@@
With increasing adoption of Modular MLLMs, research focuses on the relationship between vision encoders and MLLM performance. Tong et al. (2024b) found CLIP Radford et al. (2021) and corresponding MLLMs have similar performance trends across visual modalities, indicating CLIP features cause MLLM deficiencies in these modes, and addressed these by introducing DINOv2 Oquab et al. (2023) features. Yang et al. (2024) proposed cross-modal alignment metrics to measure vision encoder performance, fitting a binary quadratic polynomial that predicts MLLM performance using that encoder. Different from previous works, our research offers a novel perspective, demonstrating that MLLM performance correlates positively with its vision encoder’s prior knowledge.
@@ 6 Conclusion
In this paper, we introduce $Rank_{e}$ to quantify prior knowledge in vision encoder. We find that MLLM’s performance is positively correlated with prior knowledge of vision encoder, and end-to-end finetuning MLLM yields insufficient on improving low prior entity performance. To address this issue, we propose VisPRE training framework that enhances MLLM’s performance by increasing the prior knowledge within the vision encoder. Our study demonstrates a novel pathway for enhancing MLLM performance, offering substantial value for applications involving uncommon entities.
Limitations
The primary limitation of our study is the current unavailability of VQA datasets with comprehensive rare entity annotations. While our study explores MLLMs’ capabilities when confronted with uncommon entities—those inadequately represented in visual encoders’ pretraining data, most established entity-annotated datasets like S3VQA Jain et al. (2021) predominantly feature common entities. To address this challenge, we leveraged the Encyclopedia VQA Mensink et al. (2023) dataset with its diverse collection of 16.7k entity types, providing a sufficient foundation to identify and analyze less familiar entities. Nevertheless, our findings would benefit from additional specialized datasets explicitly focused on uncommon entities, which would enable a more granular analysis of visual encoders’ boundary capabilities and offer complementary insights to our current observations.
Ethics Statement
Our study utilizes MLLMs for knowledge-based VQA tasks. MLLMs may reflect biases present in the training data. Additionally, the VQA data used in our research includes pictures of landscapes and related knowledge questions, which may lead the model to generate offensive content. In this regard, we suggest users to examine the generated outputs cautiously in real-world applications.
Appendix A Datasets
Here we describe the detailed construction process of our dataset. Based on Encyclopedia-VQA Mensink et al. (2023), we constructed Knowledge-tuning and Perception-tuning datasets for each encoder-language model combination to validate Finding 2.
@@ A.1 Preprocess
Question Filtering. First, we focus on improving the parts where MLLM’s capabilities are limited by the vision encoder. Therefore, we only retained questions that could be answered by the corresponding LLM when prompted with “This is {entity_name}” instead of the actual image. Next, to ensure that there were no duplicate or similar questions for the same entity across training and test sets, we deduplicated the dataset based on (entity_name, answer) pairs. Finally, we only retained entities with three or more corresponding questions to ensure sufficient questions for dividing into training and validation sets.
Prior Calculation. We calculated $Rank_{e}$ for all entities in the filtered dataset. We examined the distribution of $Rank_{e}$ calculated using different types of encoders (CLIP Radford et al. (2021), SigLIP Zhai et al. (2023), DFN Fang et al. (2023)) across the dataset, as shown in Fig. 10. We found significant variations in $Rank_{e}$ distributions among different encoders. CLIP’s $Rank_{e}$ values were mostly concentrated in the range of $Rank_{e}<400$, with entity counts increasing as $Rank_{e}$ decreased; In contrast, SigLIP’s $Rank_{e}$ distribution is more uniform, with at least 10 entities present across most $Rank_{e}$ intervals; DFN’s $Rank_{e}$ distribution was similar to CLIP’s, with most values concentrated in the range of $Rank_{e}<400$.
Entity Sampling. For SigLIP, we divided $Rank_{e}$ into intervals of size 1000 and sampled 10 entities from each interval. For CLIP and DFN, using the same sampling strategy as SigLIP would result in insufficient sampling of entities in dense intervals, making it difficult to distinguish different levels of prior knowledge in these regions. Therefore, we adopted a sampling method that approximates the original distributions of CLIP and DFN. We sampled 10 entities from intervals of $20<Rank_{e}<=$ 2, $42<Rank_{e}<=$ 4, $84<Rank_{e}<=$ 8, …, $1024512<Rank_{e}<=$ 1024, $Rank_{e}>1024$, ensuring that the sampled distribution approximates the original distribution while retaining all entities with low prior knowledge to reflect the relationship between entity prior knowledge and model performance. Finally, we retained the questions corresponding to the sampled entities and divided the dataset into training and test sets, with statistical information shown in Table 1.
@@ A.2 Construction
For Knowledge-tuning dataset, we use the original question and answer from the Encyclopedia-VQA dataset. For Perception-tuning dataset, we replace the original question in the Knowledge-tuning dataset with cognitive question like “What is this image about?” and substitute the answers with the entity text corresponding to the image. Examples of Knowledge-tuning and Perception-tuning datasets are shown in Fig. 9.
Appendix B Evaluation Settings
We employ Llama-3.1-70B Dubey et al. (2024) to evaluate the accuracy of MLLM’s responses to VQA questions. Specifically, we provide Llama-3.1-70B with the question, entity name (wikipedia_title in prompt), ground truth answer, and MLLM’s response. The model outputs true to indicate a correct answer and false to indicate an incorrect answer. The prompt template is shown in Fig. 11, with the few_shot_examples shown in Fig. 12.
Appendix C More Cases
@@
Prompt for Llama-3.1 evaluation You are an expert evaluator tasked with assessing the correctness of model predictions. Your job is to determine if a given prediction is correct based on the provided information. Follow these strict guidelines: 1. You will be given four pieces of information: - Question: The original question asked - Wikipedia_title: The title of the Wikipedia article that corresponds to the knowledge base for the question - Answer: The correct answer(s) to the question, possibly including multiple candidates separated by "||||" - Prediction: The model’s prediction to be evaluated 2. Understand that the question is specifically about the entity described in the Wikipedia_title. 3. Compare the prediction to the answer(s), taking into account the context of the question and the Wikipedia_title. 4. Apply these strict criteria: - The prediction must be accurate and specific. - If there are multiple candidate answers separated by "||||", the prediction must match at least one of them to be considered true. - For numerical answers, the prediction must be within 10% of at least one correct answer to be considered true. - For categorical or descriptive answers, the prediction must match the key concepts or categories in at least one of the provided answers. - Partial or vague answers that don’t fully capture the specificity of any correct answer should be considered false. - Pay close attention to units, specificity, and context provided in the question, Wikipedia_title, and answer(s). 5. Your response must be exactly one word: - Output "true" if the prediction meets all the criteria for correctness. - Output "false" if the prediction fails to meet any of the criteria. 6. Do not provide any explanations or additional comments. {few_shot_examples} Remember, your task is to evaluate the correctness of the prediction based on all the information provided. Be strict in your assessment, but consider all given correct answers. Respond only with "true" or "false". Question: {question} Wikipedia_title: {wikipedia_title} Answer: {answer} Prediction: {prediction} Evaluation: